OpenStudy (anonymous):

Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2.

OpenStudy (netlopes1):

Do you know what's composite function?

OpenStudy (anonymous):

First calculate g(x)

OpenStudy (anonymous):

g(2) = -2(2) - 3 g(2) = -4 -3 = -7 g(2) = -7

OpenStudy (netlopes1):

\[f(x) = -x+4 \quad and \quad g(x)=-2x-3. \quad f(g(x))=-(-2x-3)+4 \rightarrow\]

OpenStudy (anonymous):

Now go in to calculate f(x) which would be taken from the answer from g(2) f(-7) = -(-7) + 4 f(-7) = 11 f(g(x)) = 11

OpenStudy (netlopes1):

Or...\[f(g(x))=2x+7 \rightarrow f(g(2))=2.2+7 \rightarrow f(g(2))=11. \]

OpenStudy (anonymous):

Yeah you could do whichever way is easier ! :)

OpenStudy (netlopes1):

Thanks....